OpenStudy (anonymous):

3n-5+9n=4+15n+3 Pleaseee Help :)

OpenStudy (anonymous):

12n-5-4-15n-3=0 -3n-12=0 -3n=12 n=12/-3 n=-4

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

no problem.... :)